OpenStudy (anonymous):

A cyclist rides her bike at a rate of 5 meters per second. What is this rate in kilometers per hour? How many kilometers will the cyclist travel in 3 hours?

OpenStudy (anonymous):

Rate; = km/h Distance is 3 hours = km

OpenStudy (anonymous):

@SolomonZelman @hartnn

OpenStudy (solomonzelman):

convert: 1 km = 1000 meters 1 hour = 360 seconds 5 meters per seconds= (1/200)km per 1/72 hours

OpenStudy (solomonzelman):

So far so good?

OpenStudy (anonymous):

okay yes

OpenStudy (anonymous):

18,000km/1000h

OpenStudy (solomonzelman):

How do you get this? I don't think so.

OpenStudy (anonymous):

the mins and seconds cancel out so ur left with 1h n 1km

OpenStudy (anonymous):

so makes it 18km in 1 hour

OpenStudy (solomonzelman):

1 sec / 5 meters 360 sec / 1800 meters 1 hour / 1800 meters 1 hour / 1.8 km

OpenStudy (anonymous):

so that would be the rate

OpenStudy (solomonzelman):

1 sec / 5 meters 360 sec / 1800 meters 1 hour / 1800 meters 1 hour / 1.8 km

OpenStudy (solomonzelman):

See?

OpenStudy (anonymous):

yah but the seconds cancel out

OpenStudy (solomonzelman):

what do you mean?

OpenStudy (anonymous):

5m/1s x 60s/1min x 60min/1hr x 1km/1000m

OpenStudy (anonymous):

all those cancel out besides the 1h and the 1 km

OpenStudy (anonymous):

5 x 60 x 60 = 18000

OpenStudy (anonymous):

1 x 1000 = 1000 so 18000/1000 divide it 18km/h

OpenStudy (solomonzelman):

what I did was, 1 sec / 5 meters <- both times 360 360 sec / 1800 meters 1 hour / 1800 meters <- b/c 360 sec=1 hour 1 hour / 1.8 km <- b/c 1000 km=meters, therefore 1800 meters is 1.8 km MAKES SENSE?